### v3

# Causal Deep Learning

Abstract
We derive a set of causal deep neural networks whose architectures are a consequence of tensor (multilinear) factor analysis, a framework that facilitates causal inference. Forward causal questions are addressed with a neural network architecture composed of causal capsules and a tensor transformer. Causal capsules compute a set of invariant causal factor representations, whose interactions are governed by a tensor transformation. Inverse causal questions are addressed with a neural network that implements the multilinear projection algorithm. The architecture reverses the order of the operations of a forward neural network and estimates the causes of effects. As an alternative to aggressive bottleneck dimension reduction or regularized regression that may camouflage an inherently underdetermined inverse problem, we prescribe modeling different aspects of the mechanism of data formation with piecewise tensor models whose multilinear projections produce multiple candidate solutions. Our forward and inverse questions may be addressed with shallow architectures, but for computationally scalable solutions, we derive a set of deep neural networks by taking advantage of block algebra. An interleaved kernel hierarchy results in a doubly non-linear tensor factor models. The causal neural networks that are a consequence of tensor factor analysis are data agnostic, but are illustrated with facial images. Sequential, parallel and asynchronous parallel computation strategies are described.
1 Introduction
Neural networks are being employed increasingly in high-stakes application areas, such as face recognition [70, 33, 69, 14, 93], and medical diagnosis [40, 51, 73]. Developing neural networks that offer causal explanations for correct results or failures is crucial in establishing trustworthy artificial intelligence.
Causal explanations specify the causes, the mechanism, and the conditions for replicating an observed effect [47, 24, 92]. Quantitatively, causality is the direct relationship between two events, A and B, where “A $\xrightarrow[]{\mbox{causes}}$ B” means “the effect of A is B”, a measurable and experimentally repeatable phenomena. Once verified with either experimental or observational studies, the statement “the effect of A is B” stays true regardless of new discoveries and changes in knowledge [31]. Causal explanations are a factual understanding, whereas interpretations – a concept also advocated for the development of trustworthy artificial intelligence [48] – are an understanding relative to a specific reference frame, such as a particular point of view or knowledge base. As new knowledge emerges, prior interpretations may be deemed to be inaccurate or invalid, which can undermine their reliability and usefulness in the development of trustworthy artificial intelligence. Interpretations are open to reinterpretation. Therefore, explanations should not be confused with interpretations or speculations masquerading as explanations [49, 48].
The validity and robustness of causal explanations depend on causal model specifications in conjunction with the experimental designs used for acquiring training data [63]. Generative artificial intelligence research that focuses on unsupervised deep neural networksis not well suited for drawing causal conclusions. (Determining the causal factors from unsupervised training data is an ill-posed inverse problem since different combinations of the same causal factors can result in the same outcome.) Prior unsupervised deep neural network research [21, 7, 54, 71] briefly addressed the connection to tensor (multilinear) factor analysis. Tensor (multilinear) factor analysis is a supervised learning approach that models the causal mechanisms that generate data [82, 85], estimates the effects of causes [84, 89], as well as the causes of effects, given an estimated forward model and constraints on the solution set [86, 78]. Data and its representations – such as those derived from principal or independent component analysis etc.– are factorized into a set of invariant, discriminant, and disentangled causal representations [77][(4.15)][82]. These causal representations maximize the ratio of inter- to intra-class scatter [82]. Thus, a causal tensor model is not only generative, but discriminative. In fact, the most successful deep neural network approaches are also supervised with generative and/or discriminative properties [43, 70, 53]. However, their goal has been prediction rather than causal inference.
Consequently, we derive a set of causal deep neural networks that are a consequence of causal tensor factor analysis, Figs. 1-5. Tensor factor analysis is a transparent framework for both forward [84, 82] and inverse causal inference [86, 78]. [^1] TensorFaces is a gentle introduction to causal tensor factor analysis [82, 85].
Forward causal inference is a hypothesis-driven process, as opposed to a data-driven process, that models the mechanism of data formation and estimates the effects of interventions or counterfactuals [58, 36, 68, 80]. Inverse causal “inference” estimates the causes of effects given an estimated forward model and constraints on the solution set [78, 26]. By comparison, conventional statistics and machine learning model data distributions, predict a variable co-observed with another, or perform time series forecasting.
Figure 1: A causal neural network that is an instantiation of the $M$-mode SVD [82, 85], Alg. 1. A causal neural network is composed of a set of causal capsules and a tensor transformer. Given a set of vectorized and centered training observations organized into a “data tensor” $\mathbf{\mathcal{D}}$}, causal capsules learn a set of $M$ subspaces $\{{\mathbf{U}}_{\mbox{\tiny{m}}}|1\leq m\leq M\}$ that span the $M$ causal factor representations. The tensor transformer estimates the extended core $\mathbf{\mathcal{T}}$}, which governs the interaction between causal factors. A forward causal model employs the estimated tensor decoder decoupled from the encoder to generate a new data vector. A new data vector ${\mathbf{d}}_{\mbox{\tiny{new}}}$ is the effect of new interventions, $\mbox{vec }{\mathbf{\mathcal{R}}}=(\otimes_{m=1}^{M}\hat{{\mathbf{u}}}_{m_{\mbox{\tiny{new}}}})$, where $\otimes$ is the Kronecker product. For scalable computations, each shallow autoencoder-decoder is replaced by a mathematically equivalent deep network, derived with the aid of block algebra, Fig. 3. In the upper-right corner, the causal model is summarized with graphical tensor notation [60], which may also be interpreted as a structural causal model DAG [58]. For display purposes only, the images were displayed in grid form and without being mean centered.
1.1 Causal Inference Versus Regression
Neural networks and tensor factorization methods may perform causal inference, or simply perform regression from which no causal conclusions are drawn. For causal inference, model specifications (i.e., problem definition) and experimental design for acquiring training data trump algorithmic or neural network design and analysis, Fig. 2.
Causal tensor factor analysis was employed in the analysis and recognition of facial identities [82, 80] with sparseness constraints [27, 61], as well as in analysis and recognition of facial expressions [34], human motion signatures [17, 75], and 3D sound [28]. It was employed in the rendering of arbitrary scenes, views and illuminations [84], and in the transfer of facial expressions [89] etc.Tensor factor analysis was also employed in psychometrics [74, 29, 13, 8, 44], econometrics [39, 52], chemometrics[10], and signal processing [18]. Simple tensor regression and decompositions which do not draw causal conclusions, leveraged row, column and fiber redundancies to estimate missing data [15] and to perform rank reduction [95, 90, 11, 38, 32, 6]. Recently, tensor dimensionality reduction and contractions have been employed in machine learning to reduce neural network parameters. Network parameters are organized into “data tensors”, and dimensionally reduced [46, 55, 42, 41, 57] or efficiently contracted [22].
1.2 Causal Neural Networks
Causal neural networks are composed of causal capsules and a tensor transformer, Fig. 1. Causal capsules estimate the latent variables that represent the causal factors of data formation, and a tensor transformer governs their interaction. Causal capsules may be shallow autoencoder-decoder architectures that employ linear neurons and compute a set of invariant representations  [67, 65, 64, 1, 56], as detailed in Supplemental A. The tensor transformer may be a tensor autoencoder-decoder, a shallow autoencoder-decoder whose code is the tensor product of the latent variables.
Figure 2: Same data, same algorithm, but two different model specifications (problem setups / definitions). (a) Causal Inference: The $M$-mode SVD (Alg. 1) factorizes a “data tensor” of vectorized observations into multiple subspace that span the causal factor representations. (b) Simple regression: The $M$-mode SVD factorizes a “data tensor” composed of images as a 2-way array of column or row observations into a column and row space, as well as into normalized PCA coefficients. All images are vectorized except in Fig. 2b.
Causal deep neural networks are composed by stacking autoencoders-decoders. Each autoencoder-decoder in a shallow causal neural network is replaced by mathematically equivalent deep neural network architectures that are derived by taking advantage of block algebra. An interleaved hierarchy of kernel functions [66] serves as a pre-processor that warps the data manifold for optimal tensor factor analysis. A part-based deep neural network mirrors a part-based hierarchy of tensor factor models [80, 79][77, Sec 4.4], Supplemental C. [^2] There have been a number of related transformer architectures engineered and empirically tested with success [23, 91, 50].
Inverse causal neural networks implement the multilinear projection algorithm to estimate the causes of effects [86, 78]. A neural network that addresses an underdetermined inverse problem is characterized by a wide hidden layer. Dimensionality reduction removes noise and nuisance variables [30, 72], and has the added benefit of reducing the widths of hidden layers. However, aggressive bottleneck dimensionality reduction may camouflage an inherently ill-posed problem. Alternatively or in addition to dimensionality reduction and regularized regression, we prescribe modeling different aspects of the data formation process with piecewise tensor (multilinear) models that return a set of candidate solutions [81]. Candidate solutions are gated to yield the most likely solution.
Input ${\mathbf{\mathcal{D}}}\in\mathbb{C}^{I_{0}\times\dots\times I_{\mbox{\tiny{M}}}}$, dimensions $R_{\mbox{\tiny{0}}},R_{\mbox{\tiny{1}}}\dots R_{\mbox{\tiny{m}}}\dots R_{\mbox{\tiny{M}}}$
1. Initialize ${\mathbf{U}}_{\mbox{\tiny{m}}}\mathrel{\mathop{:}}={\mathbf{I}}$, $0\leq m\leq M$
2. Iterate until convergence
  • For $m\mathrel{\mathop{:}}=0,\dots,M$,
    • 2.1
      ${\mathbf{\mathcal{X}}}\mathrel{\mathop{:}}={\mathbf{\mathcal{D}}}\times_{\mbox{\tiny{0}}}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{0}}}\times\dots\times_{\mbox{\tiny{m-1}}}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{m-1}}}\times_{\mbox{\tiny{m+1}}}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{m+1}}}\dots\times{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{M}}}$
    • 2.2
      Set ${\mathbf{U}}_{m}$ to the $\tilde{R}_{m}$ leading left-singular vectors of the SVD of ${{\mathbf{X}}}_{\mbox{\tiny{$[m]$}}}$ or SVD of ${{\mathbf{X}}}_{\mbox{\tiny{$[m]$}}}{{\mathbf{X}}{}^{\hbox{\tiny{T}}}}_{\mbox{\tiny{$[m]$}}}$].  [^3] The computation of ${\mathbf{U}}_{\mbox{\tiny{m}}}$ in the SVD ${{\mathbf{X}}}_{\mbox{\tiny{$[m]$}}}={\mathbf{U}}_{\mbox{\tiny{m}}}{\mathbf{\Sigma}}{\mathbf{V}}_{\mbox{\tiny{m}}}{}^{\hbox{\tiny{T}}}$ can be performed efficiently, depending on which dimension of ${{\mathbf{X}}}_{\mbox{\tiny{$[m]$}}}$ is smaller, by decomposing either ${{\mathbf{X}}}_{\mbox{\tiny{$[m]$}}}{{\mathbf{X}}}_{\mbox{\tiny{$[m]$}}}{}^{\hbox{\tiny{T}}}={\mathbf{U}}_{\mbox{\tiny{m}}}{\mathbf{\Sigma}}^{2}{\mathbf{U}}_{\mbox{\tiny{m}}}{}^{\hbox{\tiny{T}}}$ (note that ${\mathbf{V}}_{\mbox{\tiny{m}}}{}^{\hbox{\tiny{T}}}={\mathbf{\Sigma}}^{+}{\mathbf{U}}_{\mbox{\tiny{m}}}{}^{\hbox{\tiny{T}}}{{\mathbf{X}}}_{\mbox{\tiny{$[m]$}}}$) or by decomposing ${{\mathbf{X}}}_{\mbox{\tiny{$[m]$}}}{}^{\hbox{\tiny{T}}}{{\mathbf{X}}}_{\mbox{\tiny{$[m]$}}}={\mathbf{V}}_{\mbox{\tiny{m}}}{\mathbf{\Sigma}}^{2}{\mathbf{V}}_{\mbox{\tiny{m}}}{}^{\hbox{\tiny{T}}}$ and then computing ${\mathbf{U}}_{\mbox{\tiny{m}}}={{\mathbf{X}}}_{\mbox{\tiny{$[m]$}}}{\mathbf{V}}_{\mbox{\tiny{m}}}{\mathbf{\Sigma}}^{+}$.,  [^4] For a neural network implementation, the SVD of ${{\mathbf{X}}}_{\mbox{\tiny{$[m]$}}}$ is replaced with an autoencoder that sequentially computes the orthonormal columns of ${\mathbf{U}}_{\mbox{\tiny{m}}}$${\mathbf{V}}_{\mbox{\tiny{n}}}$ by performing gradient descent with the learning parameter $\eta$ or stochastic gradient descent [9, 62]. In Fig. 1, the autoencoders compute the columns of ${\mathbf{V}}_{\mbox{\tiny{m}}}$, where ${\mathbf{v}}_{\mbox{\tiny{m,r}}}$ is the $r^{\hbox{\small th}}$ column and it represents the weights of the $r^{\hbox{\small th}}$ neuron. Matrix ${\mathbf{V}}_{\mbox{\tiny{m,r}}}$ contains the first $r$ columns of ${\mathbf{V}}_{\mbox{\tiny{m}}}$.
        $\overbrace{\mbox{~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~}}^{\mbox{~~~~~~~~ Autoencoder~~~~~~~~}}$        For $r\mathrel{\mathop{:}}=1\dots R_{\mbox{\tiny{m}}}$.
                Iterate until convergence
      $\hskip 36.135pt\Delta{\mathbf{v}}_{\mbox{\tiny{m,r}}}\mbox{\small$(t\hskip-1.8063pt+\hskip-1.8063pt1)$}\vskip-3.61371pt$ $=$ $\eta\left({\mathbf{X}}_{\mbox{\tiny{[m]}}}-{\mathbf{V}}_{\mbox{\tiny{m,r-1}}}\mbox{\small$(t)$}{\mathbf{V}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{m,r-1}}}\mbox{\small$(t)$}{\mathbf{X}}_{\mbox{\tiny{[m]}}}\right)\underbrace{{\mathbf{X}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{[m]}}}{\mathbf{v}}_{\mbox{\tiny{m,r}}}\mbox{\small$(t)$}}_{\mbox{code}}\vskip-5.42003pt$ $\hat{{\mathbf{v}}}_{\mbox{\tiny{m,r}}}\mbox{\small$(t\hskip-1.8063pt+\hskip-1.8063pt1)$}\vskip-3.61371pt$ $=$ $\frac{\left({\mathbf{v}}_{\mbox{\tiny{m,r}}}\mbox{\small$(t)$}+\Delta{\mathbf{v}}_{\mbox{\tiny{m,r}}}\mbox{\small$(t\hskip-1.8063pt+\hskip-1.8063pt1)$}\right)}{\|{\mathbf{v}}_{\mbox{\tiny{m,r}}}\mbox{\small$(t)$}+\Delta{\mathbf{v}}_{\mbox{\tiny{m,r}}}\mbox{\small$(t\hskip-1.8063pt+\hskip-1.8063pt1)$}\|}$
3. Set ${\mathbf{\mathcal{Z}}}\mathrel{\mathop{:}}={\mathbf{\mathcal{D}}}\times_{0}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{0}}}\dots\times_{\mbox{\tiny{m}}}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{m}}}\dots\times_{\mbox{\tiny{M}}}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{M}}}\mathrel{\mathop{:}}={\mathbf{\mathcal{X}}}\times{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{M}}}$  [^5] The columns in ${{\mathbf{Z}}}_{\mbox{\tiny{$[0]$}}}$ may be computed by initializing the code of an autoencoder to $({{\mathbf{U}}}_{\mbox{\tiny{M}}}\dots\otimes{{\mathbf{U}}}_{\mbox{\tiny{m}}}\dots\otimes{{\mathbf{U}}}_{\mbox{\tiny{0}}})$, where $\otimes$ is the Kronecker product. In Fig. 1, the columns of the extended core $\mathbf{\mathcal{T}}$} are computed by initializing the code of the autoencoder with $({\mathbf{U}}_{\mbox{\tiny{M}}}\dots\otimes{\mathbf{U}}_{\mbox{\tiny{m}}}\dots\otimes{\mathbf{U}}_{\mbox{\tiny{1}}}){}^{\hbox{\tiny{T}}}$ for batch training, and $(\hat{{\mathbf{u}}}{}^{\hbox{\tiny{T}}}_{i_{\mbox{\tiny{M}}}}\dots\otimes\hat{{\mathbf{u}}}{}^{\hbox{\tiny{T}}}_{i_{\mbox{\tiny{m}}}}\dots\hat{{\mathbf{u}}}{}^{\hbox{\tiny{T}}}_{i_{\mbox{\tiny{1}}}}){}^{\hbox{\tiny{T}}}$ when training one observation, ${\mathbf{d}}_{i_{\mbox{\tiny{1}}},\dots,i_{\mbox{\tiny{M}}}}$, at a time.
Output mode matrices ${\mathbf{U}}_{0},{\mathbf{U}}_{\mbox{\tiny{1}}},\dots,{\mathbf{U}}_{\mbox{\tiny{M}}}$ and core tensor $\mathbf{\mathcal{Z}}$}.
Algorithm 1 $M$-mode SVD (parallel computation)[83, 82]
2 Forward Causal Question: “What if?”
Forward causal inference is a hypothesis-driven process that addresses the “what if” question.What if A is changed by one unit, what is the expected change in B? Causal hypotheses drive both the model specification and the experimental design for acquiring or generating training data.
2.1 Training Data
For modeling unit level effects of causes, the training data is generated by combinatorially varying each causal factor while holding the other factors fixed. The best causal evidence comes from randomized experimental studies. When randomized experiments for generating training data are unethical or infeasible, experimental studies may be approximated with carefully designed observational studies [63], such as natural experiments [2, 12, 37] or by employing the concept of transportability to transfer information from a source population where both experimental and observational studies are available to a target population where only observational studies can be conducted [59]. [^6] Gebru et al.’s “Datasheets for datasets” may aid in the approximation of experimental studies [25].
2.2 Tensor Factor Analysis Model
Within the tensor mathematical framework, an $(M+1)$-way array ${\mathbf{\mathcal{D}}}\in\mathbb{C}^{I_{\mbox{\tiny{0}}}\times I_{\mbox{\tiny{1}}}\dots\times I_{\mbox{\tiny{m}}}\dots\times I_{\mbox{\tiny{M}}}}$ is a collection of vectorized and centered observations $\{{\mathbf{d}}_{i_{\mbox{\tiny{1}}}\dots i_{\mbox{\tiny{m}}}\dots i_{\mbox{\tiny{M}}}}\in\mathbb{C}^{I_{\mbox{\tiny{0}}}}\}$ generated by $M$ causal factors. Causal factor $m$, for $m=1\dots M$, takes one of $I_{m}$ possible values, indexed by $i_{m}$, where $i_{\mbox{\tiny{m}}}=1\dots I_{m}$. An observation and an array of observations may be modeled using multilinear (tensor) principal component analysis,
${\mathbf{d}}_{i_{\mbox{\tiny{1}}},\dots,i_{\mbox{\tiny{M}}}}$ $=$ ${\mathbf{\mathcal{T}}}\times_{\mbox{\tiny{1}}}{\mathbf{\hat{u}}}{}^{\hbox{\tiny{T}}}_{i_{\mbox{\tiny{1}}}}\dots\times_{\mbox{\tiny{M}}}{\mathbf{\hat{u}}}{}^{\hbox{\tiny{T}}}_{i_{\mbox{\tiny{M}}}}+{\mathbf{e}}_{i_{\mbox{\tiny{1}}},\dots,i_{\mbox{\tiny{M}}}},$ (1)
$\ {\mathbf{\mathcal{D}}}$ $=$ ${\mathbf{\mathcal{T}}}\times_{\mbox{\tiny{1}}}{\mathbf{U}}_{\mbox{\tiny{1}}}\dots\times{\mathbf{U}}_{\mbox{\tiny{m}}}\times_{\mbox{\tiny{M}}}{\mathbf{U}}_{\mbox{\tiny{M}}}+{\mathbf{\mathcal{E}}},$ (2)
where the extended core tensor $\mathbf{\mathcal{T}}$} governs the interaction between the latent variables $\{{\mathbf{\hat{u}}}_{i_{\mbox{\tiny{m}}}}\in\mathbb{C}^{\tilde{I}_{\mbox{\tiny{m}}}}|1\leq m\leq M\mbox{ and }\tilde{I}_{\mbox{\tiny{m}}}\leq I_{\mbox{\tiny{m}}}\}$ that represent the causal factors of data formation, and $\hat{{\mathbf{u}}}{}^{\hbox{\tiny{T}}}_{i_{\mbox{\tiny{m}}}}$ is a row in ${\mathbf{U}}_{\mbox{\tiny{m}}}\in\mathbb{C}^{I_{\mbox{\tiny{m}}}\times\tilde{I}_{\mbox{\tiny{m}}}}$, Fig. 2a. [^7] A tensor $\mathbf{\mathcal{T}}$} is a multilinear mapping from a set of domain vector spaces to a range vector space, ${\mathbf{\mathcal{T}}}:\{\mathbb{C}^{I_{\mbox{\tiny{1}}}}\times\dots\times\mathbb{C}^{I_{\mbox{\tiny{M}}}}\}\xrightarrow[]{}\mathbb{C}^{I_{\mbox{\tiny{0}}}}$. In a causal tensor framework, the $M$ domain spaces span the causal factor representations and the range vector space spans the observation space. An $M$-way array of vectorized observations, $\mathbf{\mathcal{D}}$}, is sometimes informally referred to as a “data tensor”, but it is not an actual tensor. It is preferable to vectorize an image and treat it as a single observation rather than as a collection of independent column/row observations [77, App. A]. For a basic review of tensor algebra, see Supplemental B. The extended core tensor ${\mathbf{\mathcal{T}}}={\mathbf{\mathcal{D}}}\times_{\mbox{\tiny{1}}}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{1}}}\dots\times_{\mbox{\tiny{m}}}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{m}}}\dots\times_{\mbox{\tiny{M}}}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{M}}}={\mathbf{\mathcal{Z}}}\times_{\mbox{\tiny{0}}}{\mathbf{U}}_{\mbox{\tiny{0}}}$ contains the vectors that span observation subspace. [^8] In practice, the measurement mode ${\mathbf{U}}_{\mbox{\tiny{0}}}$i.e., the PCA basis matrix, and the core tensor $\mathbf{\mathcal{Z}}$} are not computed. The error ${\mathbf{e}}_{i_{\mbox{\tiny{1}}},\dots,i_{\mbox{\tiny{M}}}}\in\mathbb{C}^{I_{\mbox{\tiny{0}}}}$ is a vector in ${\mathbf{\mathcal{E}}}\in\mathbb{C}^{I_{\mbox{\tiny{0}}}\times I_{\mbox{\tiny{1}}}\dots\times I_{\mbox{\tiny{M}}}}$ and ${\mathbf{e}}_{i_{\mbox{\tiny{1}}},\dots,i_{\mbox{\tiny{M}}}}={\mathbf{\mathcal{Z}}}\times_{\mbox{\tiny{0}}}{\mathbf{E}}_{\mbox{\tiny{0}}}\times_{\mbox{\tiny{1}}}{\mathbf{\bm{\epsilon}}}{}^{\hbox{\tiny{T}}}_{i_{\mbox{\tiny{1}}}}\dots\times_{\mbox{\tiny{M}}}{\mathbf{\bm{\epsilon}}}{}^{\hbox{\tiny{T}}}_{i_{\mbox{\tiny{M}}}}$, where ${\mathbf{\bm{\epsilon}}}_{i_{\mbox{\tiny{m}}}}$ drawn from a Gaussian distribution ${\mathbf{\mathcal{N}}}({\mathbf{0}},{\mathbf{\Sigma}}_{\mbox{\tiny{m}}})$ associated with each mode $m$ ranging from 0 to M and ${\mathbf{E}}_{\mbox{\tiny{0}}}=[\dots\mathbf{\bm{\epsilon}}_{i_{\mbox{\tiny{0}}}}\dots\mathbf{\bm{\epsilon}}_{I_{\mbox{\tiny{0}}}}]{}^{\hbox{\tiny{T}}}$.
Minimizing the cost function
$L$ $=$ $\|{\mathbf{\mathcal{D}}}-{\mathbf{\mathcal{T}}}\times_{\mbox{\tiny{1}}}{\mathbf{U}}_{\mbox{\tiny{1}}}...\times_{\mbox{\tiny{m}}}{\mathbf{U}}_{\mbox{\tiny{m}}}...\times_{\mbox{\tiny{M}}}{\mathbf{U}}_{\mbox{\tiny{M}}}\|+\sum_{m=1}^{M}\lambda_{\mbox{\tiny{m}}}\|{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{m}}}\times_{\mbox{\tiny{1}}}{\mathbf{U}}_{\mbox{\tiny{m}}}-{\mathbf{I}}\|\vskip-7.22743pt$ (3)
is equivalent to maximum likelihood estimation [20] of the causal factor parameters, assuming the data was generated by the model with additive Gaussian noise. The cost
function (3) is minimized by employing $M$ alternating least squares optimizations[19, 44]
$\hskip-18.06749ptL_{\mbox{\tiny{m}}}$ $=$ $\|{\mathbf{\mathcal{X}}}_{\mbox{\tiny{m}}}-{\mathbf{\mathcal{T}}}\times_{\mbox{\tiny{m}}}{\mathbf{U}}_{\mbox{\tiny{m}}}\|+\lambda_{\mbox{\tiny{m}}}\|{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{m}}}\times_{\mbox{\tiny{1}}}{\mathbf{U}}_{\mbox{\tiny{m}}}-{\mathbf{I}}\|,$ (4)
where
$\hskip-18.06749pt{\mathbf{\mathcal{X}}}_{\mbox{\tiny{m}}}$ $\mathrel{\mathop{:}}=$ ${{\mathbf{\mathcal{D}}}\times_{\mbox{\tiny{1}}}...\times_{\mbox{\tiny{m-1}}}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{m-1}}}\times_{\mbox{\tiny{m+1}}}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{m+1}}}...\times_{\mbox{\tiny{M}}}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{M}}}}\hskip 9.03374pt\mbox{- parallel computation of all ${\mathbf{\mathcal{X}}}_{\mbox{\tiny{m}}}$\hskip-28.90755pt}$ (5)
$=$ ${\mathbf{\mathcal{X}}}_{\mbox{\tiny{m}}}\mbox{\small$(t\hskip-1.8063pt-\hskip-1.8063pt1)$}\times_{\mbox{\tiny{n}}}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{n}}}\mbox{\small$(t)$}{\mathbf{U}}_{\mbox{\tiny{n}}}\mbox{\small$(t\hskip-1.8063pt-\hskip-1.8063pt1)$},\hskip 7.22743pt\forall\mbox{\small$n\neq m$},\hskip 3.61371pt\mbox{- asynchronous parallel computation}$ (6)
$=$ $({\mathbf{\mathcal{X}}}_{\mbox{\tiny{m-1}}}\times_{\mbox{\tiny{m-1}}}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{m-1}}})\times_{\mbox{\tiny{m}}}{\mathbf{U}}_{\mbox{\tiny{m}}}\hskip 66.84996pt\mbox{- sequential computation}$ (7)
The $M$-mode SVD [82] (Alg. 1) minimizes the $M$ alternating least squares (4) in closed form by employing $M$ different SVDs. The approach is suitable for parallel (5), asynchronous (6), or sequential (7) computation.The extended core tensor $\mathbf{\mathcal{T}}$} is computed by multiplying the data tensor with the inverse mode matrices, ${\mathbf{\mathcal{T}}}={\mathbf{\mathcal{D}}}\times_{\mbox{\tiny{1}}}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{1}}}\dots\times_{\mbox{\tiny{m}}}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{m}}}\dots\times_{\mbox{\tiny{M}}}{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{M}}}$, or more efficiently as ${\mathbf{\mathcal{T}}}={\mathbf{\mathcal{X}}}_{\mbox{\tiny{m}}}\times{\mathbf{U}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{m}}}$ where ${\mathbf{\mathcal{X}}}_{\mbox{\tiny{m}}}$ is the last updated one.
2.3 Kernel Tensor Factor Analysis Model
When data $\mathbf{\mathcal{D}}$} is a tensor combination of non-linear independent causal factors, kernel multilinear independent component analysis (K-MICA) [77, Ch 4.4] employs the “kernel trick” [66, 88] as a pre-processing step which makes the data suitable for multilinear independent component analysis [85] (Alg. 2),
${\mathbf{\mathcal{D}}}$ $=$ ${\mathbf{\mathcal{T}}}\times_{\mbox{\tiny{1}}}{\mathbf{C}}_{\mbox{\tiny{1}}}\dots\times{\mathbf{C}}_{\mbox{\tiny{m}}}\dots\times_{\mbox{\tiny{M}}}{\mathbf{C}}_{\mbox{\tiny{M}}}+{\mathbf{\mathcal{E}}}\hskip 3.61371pt$ (8)
[…]
where ${\mathbf{C}}_{\mbox{\tiny{m}}}$ are the independent components, ${\mathbf{W}}_{\mbox{\tiny{m}}}$ is a rotation matrix computed either with mutual information, negentropy or higher-order cumulants, and ${\mathbf{\mathcal{E}}}={\mathbf{\mathcal{Z}}}\times_{\mbox{\tiny{0}}}{\mathbf{E}}_{\mbox{\tiny{0}}}\times_{\mbox{\tiny{1}}}{\mathbf{E}}_{\mbox{\tiny{1}}}\dots\times_{\mbox{\tiny{M}}}{\mathbf{E}}_{\mbox{\tiny{M}}}$ is the error. K-MPCA is a tensor generalization of the kernel PCA [66] and K-MICA is a tensor generalization of kernel ICA[3, 94].
To accomplish this analysis, recall that the computation of covariance matrix ${{\mathbf{D}}}_{\mbox{\tiny{$[m]$}}}{{\mathbf{D}}}_{\mbox{\tiny{$[m]$}}}{}^{\hbox{\tiny{T}}}$ involves inner products ${\mathbf{d}}{}^{\hbox{\tiny{T}}}_{i_{\mbox{\tiny{1}}}\dots i_{\mbox{\tiny{m-1}}}\,j\,i_{\mbox{\tiny{m+1}}}\dots i_{\mbox{\tiny{M}}}}{\mathbf{d}}_{i_{\mbox{\tiny{2}}}\dots i_{\mbox{\tiny{m-1}}}\,k\,i_{\mbox{\tiny{m+1}}}\dots i_{\mbox{\tiny{M}}}}$ between pairs of data points in the data tensor $\mathbf{\mathcal{D}}$} associated with causal factor $m$, for $m=1,\dots,M$ (Step 2.2 in Algorithm 1). We replace the inner products with a generalized distance measure between images, $K({\mathbf{d}}_{i_{1}\dots i_{m-1}\,j\,i_{m+1}\dots i_{\mbox{\tiny{M}}}},{\mathbf{d}}_{i_{2}\dots i_{m-1}\,k\,i_{m+1}\dots i_{\mbox{\tiny{M}}}})$, where $K(\cdot,\cdot)$ is a suitable kernel function (Table 1) that corresponds to an inner product in some expanded feature space. This generalization naturally leads us to a Kernel Multilinear PCA (K-MPCA) Algorithm, where the covariance computation is replaced by
$\left[{{\mathbf{D}}}_{\mbox{\tiny{$[m]$}}}{{\mathbf{D}}}_{\mbox{\tiny{$[m]$}}}{}^{\hbox{\tiny{T}}}\right]_{jk}$ $\mathrel{\mathop{:}}=$ $\sum_{i_{1}=1}^{I_{1}}\dots\sum_{i_{m-1}=1}^{I_{m-1}}~\sum_{i_{m+1}=1}^{I_{m+1}}\dots\sum_{i_{\mbox{\tiny{M}}}=1}^{I_{\mbox{\tiny{M}}}}$ (10)
$K({\mathbf{d}}_{i_{1}\cdots i_{m-1}\,j\,i_{m+1}\cdots i_{\mbox{\tiny{M}}}},{\mathbf{d}}_{i_{1}\cdots i_{m-1}\,k\,i_{m+1}\cdots i_{\mbox{\tiny{M}}}}).$
When a causal factor is a combination of multiple independent sources that are causal in nature, we employ a rotation matrix $\mathbf{W}$} to identify them. The rotation matrix is computed by employing either mutual information, negentropy, or higher-order cumulants [16, 5, 35, 4]. A Kernel Multilinear ICA (K-MICA) Algorithm is a kernel generalization of the multilinear independent component analysis (MICA) algorithm [85]. Algorithm 2 simultaneously specifies both K-MPCA and K-MICA algorithms. A scalable tensor factor analysis represents an observation as a hierarchy of parts and wholes [80, 79].
Input the data tensor ${\mathbf{\mathcal{D}}}\in\mathbb{C}^{I_{0}\times\dots\times I_{M}}$, where mode $m=0$ is the measurement mode, and the desired ranks are $\tilde{R}_{1},\dots,\tilde{R}_{M}$.
Initialize ${\mathbf{C}}_{\mbox{\tiny{m}}}={\mathbf{I}}$, $\forall 0\leq m\leq M$
Iterate until convergence.
  1. 1.
    For $m\mathrel{\mathop{:}}=1,\dots,M$
    1. (a)
      Set ${\mathbf{\mathcal{X}}}_{\mbox{\tiny{m}}}\mathrel{\mathop{:}}={\mathbf{\mathcal{D}}}\times_{1}{\mathbf{C}}_{1}^{+}\dots\times_{m-1}{\mathbf{C}}_{m-1}^{+}\times_{m+1}{\mathbf{C}}_{m+1}^{+}\dots\times_{M}{\mathbf{C}}_{M}^{+}.$
    2. (b)
      Compute the elements of the mode-$m$ covariance matrix using kernel functions, Table1, for $j,k\mathrel{\mathop{:}}=1,\dots,I_{m}$:
      $\left[{{\mathbf{X}}_{\mbox{\tiny{m}}}}_{\mbox{\tiny{$[m]$}}}{{\mathbf{X}}}_{\mbox{\tiny{$[m]$}}}{}^{\hbox{\tiny{T}}}\right]_{\mbox{\tiny{jk}}}\thinspace\mathrel{\mathop{:}}=\hskip-3.61371pt\sum_{i_{1}=1}^{I_{1}}...\thinspace\sum_{i_{\mbox{\tiny{m-1}}}=1}^{I_{m-1}}~\sum_{i_{\mbox{\tiny{m+1}}}=1}^{I_{\mbox{\tiny{m+1}}}}...\sum_{i_{\mbox{\tiny{M=1}}}}^{I_{M}}\thinspace K({\mathbf{x}}_{i_{\mbox{\tiny{1}}}...i_{\mbox{\tiny{m-1}}}\,j\,i_{\mbox{\tiny{m+1}}}...i_{\mbox{\tiny{M}}}},{\mathbf{x}}_{i_{\mbox{\tiny{1}}}...i_{\mbox{\tiny{m-1}}}\,k\,i_{\mbox{\tiny{m+1}}}...i_{\mbox{\tiny{M}}}}).$ (11)
    3. (c)  [^9] Every SVD step may be autoencoder-decoder. See Algorithm 1, Footnotes 3 and 4.
      See Fig. 3 for a scalable neural network implementation.
[…]
  2. 2.
[…]
Output the converged extended core tensor ${\mathbf{\mathcal{T}}}\in\mathbb{C}^{I_{0}\times\tilde{R}_{1}\times\dots\times\tilde{R}_{M}}$ and causal factor mode matrices ${\mathbf{C}}_{1},\dots,{\mathbf{C}}_{M}$.
Algorithm 2 Kernel Tensor Factor Analysis [77, Sec 4.4],[85]
Kernel Multilinear Independent Component Analysis (K-MICA) and
Kernel Principal Component Analysis (K-MPCA).
Linear kernel: $K({\mathbf{u}},{\mathbf{v}})={\mathbf{u}}{}^{\hbox{\tiny{T}}}{\mathbf{v}}={\mathbf{u}}\cdot{\mathbf{v}}$ Polynomial kernel of degree $d$: $K({\mathbf{u}},{\mathbf{v}})=({\mathbf{u}}{}^{\hbox{\tiny{T}}}{\mathbf{v}})^{d}$ Polynomial kernel up to degree $d$: $K({\mathbf{u}},{\mathbf{v}})=({\mathbf{u}}{}^{\hbox{\tiny{T}}}{\mathbf{v}}+1)^{d}$ Sigmoidal kernel: $K({\mathbf{u}},{\mathbf{v}})=\tanh(\alpha\,{\mathbf{u}}{}^{\hbox{\tiny{T}}}{\mathbf{v}}+\beta)$ Gaussian (radial basis function (RBF)) kernel: $K({\mathbf{u}},{\mathbf{v}})=\exp\left(-{\|{\mathbf{u}}-{\mathbf{v}}\|^{2}\over 2\sigma^{2}}\right)$
Table 1: Common kernel functions. Kernel functions are symmetric, positive semi-definite functions corresponding to symmetric, positive semi-definite Gram matrices. The linear kernel does not modify or warp the feature space.
2.4 Neural Network Architecture
[…]
${\mathbf{d}}_{i_{\mbox{\tiny{1}}},\dots,i_{\mbox{\tiny{m}}}\dots i_{\mbox{\tiny{M}}}}$ $=$ ${\mathbf{\mathcal{T}}}\times_{\mbox{\tiny{1}}}\hat{{\mathbf{u}}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{1}}}\times\dots\times_{\mbox{\tiny{m}}}\hat{{\mathbf{u}}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{m}}}\times\dots\times_{\mbox{\tiny{M}}}\hat{{\mathbf{u}}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{M}}}$ (12)
$=$ ${{\mathbf{T}}}_{\mbox{\tiny{$[0]$}}}({\mathbf{u}}{}^{\hbox{\tiny{T}}}_{i_{\mbox{\tiny{M}}}}\otimes\dots\otimes{\mathbf{u}}{}^{\hbox{\tiny{T}}}_{i_{\mbox{\tiny{m}}}}\dots\otimes{\mathbf{u}}{}^{\hbox{\tiny{T}}}_{i_{\mbox{\tiny{1}}}}){}^{\hbox{\tiny{T}}}$. (13)
To address a set of arbitrarily non-linear causal factors, each autoencoder employs kernel functions (Algorithm 2,Table 1).
2.5 Causal Deep Networks and Scalable Tensor Factor Analysis:
      (b)      (c) (d) (e)
Refer to caption Refer to caption
              (f)                (g)
Figure 3: Deep neural network. Subfigures (a-e) depict (3140) and  [77, pg.38-40].
(a) The mode matrix computation ${\mathbf{U}}_{\mbox{\tiny{m}}}$ may be thought as a constrained cluster-based PCA that is rewritten in terms of block SVDs. Matrix4izing $\mathbf{\mathcal{D}}$} may be viewed as a concatenation of “cluster” data. The matrix $\mathbf{W}$} transforms the basis matrix ${\mathbf{V}}^{(n)}_{\mbox{\tiny{0}}}$ such that the causal factor representation ${\mathbf{U}}_{\mbox{\tiny{m}}}$ is the same regardless of cluster membership. In a tensor model, there are $M$ different constrained cluster-based PCAs. (b) Mode matrix ${\mathbf{U}}_{\mbox{\tiny{m}}}$ computation using a single autoencoder-decoder. (c) Mode matrix computation as a hierarchy of autoencoder-decoders (i.e., feed-forward [45]), (d) Mode matrix computation written as a deep learning model (e) Concurrent-autoencoders; i.e., constrained cluster-based autoencoders. (f) Forward causal model with a set of capsules implemented by deep neural networks. For a parallel, synchronized or asynchronous computation, we break the chain links and shuttle causal information, ${\mathbf{U}}_{\mbox{\tiny{m}}}$, between capsules to compute ${\mathbf{\mathcal{X}}}(t+1)$ for the next iteration. (g) Each capsule in (f) may be replaced with a part-based deep neural network by permuting the rows in ${{\mathbf{D}}{}^{\hbox{\tiny{T}}}}_{\mbox{\tiny{$[m]$}}}$ with $\mathbf{P}$}, segmented by ${\mathbf{H}}_{\mbox{\tiny{m}}}$, which is efficiently trained with a part-based hierarchy of autoencoders, Fig. 4.
For a scalable architecture, we leverage the properties of block algebra. Shallow autoencoders are replaced with either a mathematically equivalent deep neural network that is a part-based hierarchy of autoencoders-decoders (i.e., feed-forward of Restricted Boltzman Machines [45]), or a set of concurrent autoencoders-decoders, Fig. 3.
For example, the orthonormal subspace of a data batch, ${\mathbf{D}}\in\mathbb{C}^{I_{\mbox{\tiny{0}}}\times I_{\mbox{\tiny{1}}}}$ that has $I_{\mbox{\tiny{0}}}$ measurements and $I_{\mbox{\tiny{1}}}$ observations may be computed by recursively subdividing the data, analyzing the data blocks, dimensionally reducing their representations, and finally merging the information
$\hskip-3.61371pt{\mathbf{D}}$ $=$ $\left[\begin{array}[]{cc}{\mathbf{D}}_{\mbox{\tiny{A}}}\\ {\mathbf{D}}_{\mbox{\tiny{B}}}\end{array}\right]=\left[\begin{array}[]{c}{\mathbf{U}}_{\mbox{\tiny{A}}}{\mathbf{S}}_{\mbox{\tiny{A}}}{\mathbf{V}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{A}}}\\ {\mathbf{U}}_{\mbox{\tiny{B}}}{\mathbf{S}}_{\mbox{\tiny{B}}}{\mathbf{V}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{B}}}\end{array}\right]=\left[\begin{array}[]{cc}{\mathbf{U}}_{\mbox{\tiny{A}}}&{\mathbf{0}}\\ {\mathbf{0}}&{\mathbf{U}}_{\mbox{\tiny{B}}}\end{array}\right]\underbrace{\left[\begin{array}[]{c}{\mathbf{S}}_{\mbox{\tiny{A}}}{{\mathbf{V}}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{A}}}\\ {\mathbf{S}}_{\mbox{\tiny{B}}}{{\mathbf{V}}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{B}}}\end{array}\right]}_{\mbox{SVD}}=\left[\begin{array}[]{cc}{\mathbf{U}}_{\mbox{\tiny{A}}}&{\mathbf{0}}\\ {\mathbf{0}}&{\mathbf{U}}_{\mbox{\tiny{B}}}\end{array}\right]{\mathbf{W}}{\mathbf{\Sigma}}{\mathbf{V}}{}^{\hbox{\tiny{T}}}$ (24)
$=$ $\left[\begin{array}[]{cc}{\mathbf{U}}_{\mbox{\tiny{A}}}&{\mathbf{0}}\\ {\mathbf{0}}&{\mathbf{U}}_{\mbox{\tiny{B}}}\end{array}\right]\left[\begin{array}[]{cc}{\mathbf{W}}_{\mbox{\tiny{AA}}}&{\mathbf{W}}_{\mbox{\tiny{AB}}}\\ {\mathbf{W}}_{\mbox{\tiny{BA}}}&{\mathbf{W}}_{\mbox{\tiny{BB}}}\end{array}\right]{\mathbf{\Sigma}}{\mathbf{V}}{}^{\hbox{\tiny{T}}}=\left[\begin{array}[]{c}{\mathbf{U}}_{\mbox{\tiny{A}}}{\mathbf{W}}_{\mbox{\tiny{AA}}}\\ {\mathbf{U}}_{\mbox{\tiny{B}}}{\mathbf{W}}_{\mbox{\tiny{BB}}}\end{array}\right]{\mathbf{\Sigma}}{\mathbf{V}}{}^{\hbox{\tiny{T}}}={\mathbf{U}}{\mathbf{\Sigma}}{\mathbf{V}}{}^{\hbox{\tiny{T}}}$. (31)
The rotation matrix $\mathbf{W}$} transforms basis matrices, ${\mathbf{U}}_{\mbox{\tiny{A}}}$ and ${\mathbf{U}}_{\mbox{\tiny{B}}}$, which span the subspaces of the observations in the data blocks ${\mathbf{D}}_{\mbox{\tiny{A}}}$ and ${\mathbf{D}}_{\mbox{\tiny{B}}}$, such that the observations in ${\mathbf{D}}_{\mbox{\tiny{A}}}$ and the corresponding observations in ${\mathbf{D}}_{\mbox{\tiny{B}}}$ no longer have distinct representations ${\mathbf{V}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{A}}}$ and ${\mathbf{V}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{B}}}$, but have the same representations ${\mathbf{V}}{}^{\hbox{\tiny{T}}}$. The representations ${\mathbf{V}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{A}}}$ and ${\mathbf{V}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{B}}}$ may be dimensionally reduced before moving onto the next step. [^10] Block algebra may be employed if the tensor model is multilinear (tensor) principal component analysis (MPCA), multilinear (tensor) independent component analysis (MICA)[85], Kernel-MPCA or Kernel-MICA[77].
Computing causal factor representations, the mode matrices ${\mathbf{U}}_{\mbox{\tiny{m}}}$ of an MPCA tensor model, is equivalent to computing a $M$ different of mutually constrained, cluster-based PCA,  [77, pg.38-40], Fig. 3a. When dealing with data that can be separated into clusters, the standard machine learning approach is to compute a separate PCA. When data from different clusters are generated by the same underlying process (e.g., facial images of the same people under different viewing conditions), the data blocks can be concatenated in the measurement mode and the common causal factor can be modeled by one PCA. However, for a scalable solution, we employ block algebra (31) and compute a set of constrained cluster-based PCAs, i.e., a set of concurrent PCAs. Thus, we define a constrained, cluster-based PCA as the computation of a set of PCA basis vectors, such that the latent representation is constrained to be the invariant of the cluster membership.
Refer to caption       Refer to caption
   (a)      (b)
Figure 4: (a) Causal capsules may be implemented with a part-based hierarchy of autoencoders. The dataset is permuted by $\mathbf{P}$}, segmented and filtered by ${\mathbf{H}}_{\mbox{\tiny{m}}}$ that is mode dependent. (b) Implementing the capsules with a part-based hierarchy of autoencoders is equivalent to performing M-mode Block SVD[80, Sec IV][76].
In the context of our multifactor data analysis, we define a cluster as a set of observations for which all factors are fixed but one. For every tensor mode, there are $N_{\mbox{\tiny{m}}}=I_{1}I_{2}\dots I_{m-1}I_{m+1}\dots I_{\mbox{\tiny{M}}}$ possible clusters and the data in each cluster varies with the same causal mode. The constrained, cluster-based PCA concatenates the clusters in the measurement mode and analyzes the data with a linear model, such as PCA.
To see this, let ${\mathbf{\mathcal{D}}}_{i_{1}\dots i_{m-1}i_{m+1}\dots i_{\mbox{\tiny{M}}}}\in\mathbb{C}^{I_{0}\times 1\times 1\dots\times 1\times I_{\mbox{\tiny{m}}}\times 1\dots\times 1}$ denote a subtensor of $\mathbf{\mathcal{D}}$} that is obtained by fixing all causal factor modes but mode $m$ and mode 0 (the measurement mode). Matrixizing this subtensor in the measurement mode we obtain ${{\mathbf{D}}_{i_{1}\dots i_{m-1}i_{m+1}\dots i_{\mbox{\tiny{M}}}}}_{\mbox{\tiny{$[0]$}}}\in\mathbb{C}^{I_{0}\times I_{\mbox{\tiny{m}}}}$. This data matrix comprises a cluster of data obtained by varying causal factor $m$, to which one can traditionally apply PCA. Since there are $N_{\mbox{\tiny{m}}}=I_{1}I_{2}\dots I_{m-1}I_{m+1}\dots I_{\mbox{\tiny{M}}}$ possible clusters that share the same underlying space associated with factor $m$, the data can be concatenated and PCA performed in order to extract the same representation for factor $m$ regardless of the cluster. Now, consider the MPCA computation of mode matrix ${\mathbf{U}}_{\mbox{\tiny{m}}}$, Fig. 3a, which can be written in terms of matrixized subtensors as
[…]
This is equivalent to computing a set of $N_{\mbox{\tiny{m}}}=I_{1}I_{2}\dots I_{m-1}I_{m+1}\dots I_{\mbox{\tiny{M}}}$ cluster-based PCAs concurrently by combining them into a single statistical model and representing the underlying causal factor $m$ common to the clusters. Thus, rather than computing a separate linear PCA model for each cluster, MPCA concatenates the clusters into a single statistical model and computes a representation (coefficient vector) for mode $m$ that is invariant relative to the other causal factor modes 1,,(m1),(m+1),,M1,...,(m-1),(m+1),...,M. For a scalable solution, we rotate the cluster-based PCA basis vectors, such that the data blocks have the same representation regardless of cluster memebership. Thus, MPCA performs multiple constrained, cluster-based PCA. To clarify the relationship, let us number each of the matrices ${{\mathbf{D}}_{i_{1}\dots i_{m-1}i_{m+1}\dots i_{\mbox{\tiny{M}}}}}_{\mbox{\tiny{$[m]$}}}={\mathbf{D}}_{\mbox{\tiny{m}}}^{(n)}$ with a parenthetical superscript $1\leq n=1+\sum_{k=1,k\neq m}^{M}(i_{n}-1)\prod_{l=1,l\neq m}^{k-1}I_{l}\leq N_{\mbox{\tiny{m}}}$. Let each of the cluster SVDs be ${\mathbf{D}}_{\mbox{\tiny{m}}}^{(n)}={\mathbf{U}}_{\mbox{\tiny{m}}}^{(n)}{\mathbf{\Sigma}}_{\mbox{\tiny{m}}}^{(n)}{{\mathbf{V}}_{m}^{(n)}}{}^{\hbox{\tiny{T}}}$, and
${{\mathbf{D}}}_{\mbox{\tiny{$[m]$}}}$ $=$ $\underbrace{\left[\begin{array}[]{c}{\mathbf{U}}_{\mbox{\tiny{m}}}^{(1)}{\mathbf{\Sigma}}_{\mbox{\tiny{m}}}^{(1)}\dots\mathbf{U}_{\mbox{\tiny{m}}}^{(N_{\mbox{\tiny{M}}})}{\mathbf{\Sigma}}_{\mbox{\tiny{m}}}^{(N_{\mbox{\tiny{m}}})}\end{array}\right]}_{\mbox{SVD}}\mathop{\mathrm{diag}}([\begin{array}[]{c}{\mathbf{V}}^{(1)}_{\mbox{\tiny{m}}}\dots\mathbf{V}^{(N_{\mbox{\tiny{m}}})}_{\mbox{\tiny{m}}}\end{array}]){}^{\hbox{\tiny{T}}}$ (35)
$=$ ${\mathbf{U}}_{\mbox{\tiny{m}}}{\mathbf{\Sigma}}_{\mbox{\tiny{m}}}{\mathbf{W}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{m}}}\mathop{\mathrm{diag}}([\begin{array}[]{ccc}{\mathbf{V}}_{\mbox{\tiny{m}}}^{(1)}&\dots&{\mathbf{V}}_{\mbox{\tiny{m}}}^{(N_{\mbox{\tiny{m}}})}\end{array}]){}^{\hbox{\tiny{T}}}$, (37)
$=$ ${\mathbf{U}}_{\mbox{\tiny{m}}}{\mathbf{\Sigma}}_{\mbox{\tiny{m}}}[\begin{array}[]{ccc}{\mathbf{V}}^{(1)}_{\mbox{\tiny{m}}}{\mathbf{W}}^{(1)}_{\mbox{\tiny{m}}}&\dots&{\mathbf{V}}^{(N_{\mbox{\tiny{m}}})}_{\mbox{\tiny{m}}}{\mathbf{W}}^{(N_{\mbox{\tiny{m}}})}_{\mbox{\tiny{m}}}\end{array}]{}^{\hbox{\tiny{T}}}$ (39)
$=$ ${\mathbf{U}}_{\mbox{\tiny{m}}}{\mathbf{\Sigma}}_{\mbox{\tiny{m}}}{\mathbf{V}}_{\mbox{\tiny{m}}}{}^{\hbox{\tiny{T}}}$, (40)
where $\mathop{\mathrm{diag}}(\cdot)$ denotes a diagonal matrix whose elements are each of the elements of its vector argument. The mode matrix ${\mathbf{V}}^{(n_{\mbox{\tiny{m}}})}_{\mbox{\tiny{m}}}$ is the measurement matrix ${\mathbf{U}}^{(n_{\mbox{\tiny{m}}})}_{\mbox{\tiny{0}}}$ (${\mathbf{U}}^{(n_{\mbox{\tiny{m}}})}_{\mbox{\tiny{x}}}$ when the measurements are image pixels) that contains the eigenvectors spanning the observed data in cluster $n_{\mbox{\tiny{m}}}$, $1\leq n_{\mbox{\tiny{m}}}\leq N_{\mbox{\tiny{m}}}$. MPCA can be thought as computing a rotation matrix, ${\mathbf{W}}_{\mbox{\tiny{m}}}$, that contains a set of blocks ${\mathbf{W}}^{(n)}_{\mbox{\tiny{m}}}$ along the diagonal that transform the PCA cluster eigenvectors ${\mathbf{V}}^{(n_{\mbox{\tiny{m}}})}_{\mbox{\tiny{m}}}$ such that the mode matrix ${\mathbf{U}}_{\mbox{\tiny{m}}}$ is the same regardless of cluster membership (3540), Fig 3. The constrained “cluster”-based PCAs may also be implemented with a set of concurrent “cluster”-based PCAs, Fig. 3e.
Causal factors of object wholes may be computed efficiently from their parts, by applying a permutation matrix $\mathbf{P}$} and creating part-based data clusters with a segmentation filter ${\mathbf{H}}_{\mbox{\tiny{m}}}$, where ${\mathbf{\mathcal{D}}}\hskip 3.61371pt\,\times^{\thinspace\hbox{\tiny{}T}}_{\mbox{\tiny{m}}}{\mathbf{H}}_{\mbox{\tiny{m}}}{\mathbf{P}}\Leftrightarrow{\mathbf{H}}_{\mbox{\tiny{m}}}{\mathbf{P}}{{{\mathbf{D}}}_{\mbox{\tiny{$[m]$}}}}{}^{\hbox{\tiny{T}}}$. A deep neural network can be efficiently trained with a hierarchy of part-based autoencoders, Fig. 4. A computation that employs a part-based hierarchy of autoencoders parallels the Incremental M-mode Block SVD [80, 76, Sec. IV]. A data tensor is recursively subdivided into data blocks, analyzed in a bottom-up fashion, and the results merged as one moves through the hierarchy. The computational cost is the cost of training one autoencoder, $\mathcal{O}(T)$, times $\mathcal{O}(logN_{\mbox{\tiny{M}}})$, the total number of autoencoders trained for each factor matrix, $\mathcal{O}(TlogN_{\mbox{\tiny{m}}})$. If the causal neural network is trained sequentially, the training cost for one time iteration is $\mathcal{O}(MTlog\bar{N})$, where $\bar{N}$ is the average number of clusters across the $M$ modes.
3 Inverse Causal Question: “Why?”
Inverse causal inference addresses the “why” question and estimates the causes of effects given an estimated forward causal model and a set of constraints that reduce the solution set [^11] As discussed earlier in the paper, different combinations of the same causal factors can lead to the same outcome. In imaging, this phenomena may result in visual illusions. This is a many-to-one problem, and its inverse is ill posed without constraints. and render the problem well-posed [26, 78, 86].
Multilinear tensor factor analysis constrains causal factor representations to be unitary vectors. Multilinear projection [86, 78] relies on this constraint and performs multiple regularized regressions. One or more unlabeled test observations that are not part of the training data set are multilinearly projected into the causal factor spaces,
[…]
Neural architectures addressing underdetermined inverse problems are characterized by hidden layers that are wider than the input layer; i.e., the dimensionality of $vec({\mathbf{\mathcal{R}}})$ is larger than the number of measurements in $\mathbf{d}$}. Dimensionality reduction reduces noise, and the width of the hidden layers [30]. However, it can also camouflage an inherently underdetermined inverse problem. Adding sparsity, non-negativity constraints[87], etc., can further reduce the solution set in a principled way. Alternatively or in addition, one can determine a set of candidate solutions by modeling different aspects of the mechanism of data formation as piecewise tensor (multilinear) factor models. A single multilinear projection [86, 78] is replaced with multiple multilinear projections. Vasilescu and Terzopoulos [81] rewrote the forward multilinear model in terms of multiple piecewise linear models that were employed to perform multiple linear projections and produced multiple candidate solutions that were gated to return the most likely solution.
[…]
4 Conclusion
We derive a set of shallow and deep causal neural networks that are a consequence of causal tensor factor analysis. Causal neural networks are composed of causal capsules and a tensor transformer. Causal capsules compute invariant causal factor representations, whose interaction are governed by a tensor transformation. An inverse causal neural network implements the multilinear projection and estimates the causes of effects. As an alternative to aggressive “bottleneck” dimensionality reduction that may camouflage an inherently underdetermined inverse problem, the mechanism of data formation is modeled as piecewise tensor (multilinear) models, and inverse causal neural networks perform multiple multilinear projections that result in multiple candidate solutions, which may be gated to yield the most likely solution.
Acknowledgement: Demetri Terzopoulos and Ernest Davis provided invaluable feedback on various drafts of this paper.
Supplemental
A. PCA computation with a Hebb Autoencoder
B. Relevant Tensor Algebra
C. Compositional Hierarchical Block TensorFaces